# Multicritical quantum sensors driven by symmetry-breaking

### v1

@@ Abstract
Quantum criticality has been demonstrated as a useful quantum resource for parameter estimation. This includes second-order, topological and localization transitions. In all these works reported so far, gap-to-gapless transition at criticality has been identified as the ultimate resource for achieving the quantum enhanced sensing, although there are several important concepts associated with criticality, such as long-range correlation, symmetry breaking. In this work, we analytically demonstrate that symmetry-breaking can drive a quantum enhanced sensing in single- or multiparameter estimation. We show this in the well-known Kitaev model, a lattice version of the 1D $p$-wave superconductor, which consists of a pairing term and an onsite potential term. The model is characterized by two critical lines and a multi-critical point at the intersection of these two lines. We show that Heisenberg scaling can be obtained in precision measurement of the superconducting coupling by preparing the system at or near the multicritical point despite the fact that parameter variation follows the critical lines, i.e., without an explicit requirement of gap-to-gapless transition. Quantum enhancement in such situations solely occurs due to a global $U(1)$ symmetry-breaking by the pairing term. Extending our analysis in the realm of multiparameter estimation we show that it is possible to obtain super-Heisenberg scaling by combining the effects of symmetry-breaking and gapless-to-gapped transition.
@@ I Introduction
The quantum Cramér-Rao bound provides the ultimate limit of precision within which an unknown parameter can be estimated in the quantum systems [1, 2, 3, 4]. The uncertainty $\epsilon_{q}$ in parameter estimation is related to a very special quantity, quantum Fisher information. The bound suggests that $\epsilon_{q}\geq(MF_{Q})^{-1}$ for the quantum systems, where $F_{Q}$ is the quantum Fisher information (QFI) and $M$ is the number of repetition of the sensing protocol. Quantum systems might provide a true quantum edge over classical systems when it comes to accurately measuring an unknown variable. This is demonstrated by how the uncertainty scales with the size of the probe. Given $\epsilon_{q}\propto L^{-\nu}$, where $L$ is the prob size and $\nu$ is the associate scaling exponent, the classical probes can at-best scale linearly with system size, i.e., $\nu=1$, which corresponds to the standard quantum limit (SQL). However, it is possible to exploit quantum phenomena, such as quantum correlation or cooperative phenomena, for achieving quantum-enhanced sensitivity, where $\nu>1$. The so-called Heisenberg limit (HL) corresponds to $\nu=2$ [5]. Subsequently, quantum sensing [6, 2, 7] has taken the centre stage of quantum technology with several applications, and has been investigated from different points of views [5, 8, 9, 10, 11, 12].
@@
The cooperative quantum phenomena, such as quantum phase transitions in quantum many-body (QMB) systems, has turned out to be an extremely useful quantum resource for engineering new types of quantum sensors. Quantum criticality, in its various forms, have been utilized to attain HL limit, or for even realizing beyond HL limits. This includes symmetry-breaking second-order quantum phase transitions [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 17, 34], as well as symmetry-protected quantum phase transitions, such as localization transition [35, 36] and topological phase transition (TPT) [37, 37, 38, 39, 40, 41, 42], and non-Hermitian topological systems [43, 44, 45, 46]. There are two different key motivations for probing different kinds of quantum criticality - first, it is important to understand the basic ingredients leading to quantum-enhanced sensitivity, e.g. the role of gap closing or symmetry-breaking; secondly, there is a constant quest for identifying experimentally realizable QMB systems with super-Heisenberg scaling property characterized by large scaling exponent. It has been shown that the gap closing at criticality leads to a scaling of $F_{q}\sim L^{(2/\nu)}$, where $L$ is the system size, and the critical exponent $\nu$ characterizing the divergence of the correlation length (localization length) near the criticality in case of the second-order quantum phase transition (localization transition). Although there are many key aspects associated with quantum criticality, such as symmetry breaking, gap closing, long-range correlation, only the gap closing has been identified as the crucial ingredient for obtaining quantum enhanced sensitivity, implying $\nu<2$.
Figure 1: Schematic for quantum enhanced sensing assisted by symmetry-breaking: The phase diagram shows different topological phases in the $\mu$-$\Delta$ plane. The red and the green lines are gapless. In the left panel, the protocol of sensing $\Delta$ is presented. $\mathcal{F}_{\Delta\Delta}$ has an quantum enhanced scaling when the system is prepared at the critical line of $\mu=\mu_{c}$ (near the multicritical point, at A), and driven to, say B. We note that this is enhancement of scaling is not because of gap-to-gapless transition, but rather due to breaking of $U(1)$ symmetry. In the multi-parameter regime, one can envision a case where system is prepared at C, and both the parameters are maneuvered along the directed arrows. This leads to a quantum enhanced scaling for the multi-parameter precision estimator, $\mathcal{G}\sim L^{6}$. This involves, both, a gap-to-gapless transition and the symmetry-breaking.
In this work we show that gap-to-gapless transition is always not required for quantum-enhanced sensing. Symmetry-breaking at multi-criticality can also be sufficient for it. In our knowledge such symmetry-breaking quantum advantage for quantum sensing has not been reported in literature previously. We consider well-known Kitaev model on a one-dimensional lattice with nearest neighbor hopping, onsite potential with strength $\mu$ and a $p$-wave pairing term with amplitude $\Delta$, for demonstrating the idea analytically. The model has two critical lines at $\Delta=0$ and $|\mu|=2$, crossing whom one observes discrete jumps in a topological invariant. It hosts multi-critical points at the intersections of these two critical lines. The system is prepared at or near the multi-critical point. In the context of single parameter estimation we show that precision estimation of $\Delta$ is still possible, even though then parameter variation in $\Delta$ is restricted to follow the critical line at $\mu=2$, where it does not encounter any gapless-to-gap transition. We show that the system still can achieve Heisenberg scaling. We expand our interest in the realm of multi-parameter sensing, where we perform simultaneous estimation of both, $\mu$ and $\Delta$. Despite the fact that parameter variations follow the gapless critical lines, super-Heisenberg scaling is observed for multi-parameter estimation. In both the cases the quantum enhanced sensing is essentially facilitated via a symmetry-breaking at $\Delta=0$, at which the system has a $U(1)$ symmetry. This forms the central result presented in this work.
Figure 2: The variation of $\mathcal{F}_{\mu\mu}$ around the multicriticality : (a) $\mathcal{F}_{\mu\mu}$ for system sizes N = 400 (solid blue line), 600 (dot-dashed orange line), 800 (dotted green line), 1000 (dashed red line) are presented against $\mu$, where in the x-axis $\mu_{c}=2$ has been subtracted from $\mu$. $\Delta=0.001$ is set for all cases. This is a representative case, we observe a similar nature for various other system-sizes. (b) The finite-size scaling of $\mathcal{F}_{\mu\mu}$ is presented at $\mu=2$ and two different $\Delta$. Considering $\Delta=10^{-7}$, we observe $\mathcal{F}_{\mu\mu}^{a}\sim L^{6.0}$ (circular markers) and $\Delta=0.7$, we observe $\mathcal{F}_{\mu\mu}^{b}\sim L^{2.0}$ (square markers). The straight lines are the best fit. (c) We present the scaling exponents $\beta$ of $\mathcal{F}_{\mu\mu}$ against the parameter $\Delta$. We identify three regions, (i) low $\Delta$ regime: $\Delta<10^{-3}$ represented by light blue color where $\mathcal{F}_{\mu\mu}\sim L^{6}$; (ii) intermediate $\Delta$ regime: $10^{-3}\leq\Delta\leq 10^{-2}$ is represented by light green color, where $\mathcal{F}_{\mu\mu}\sim L^{\beta}$, $\beta\in(2,6)$; (iii) large $\Delta$ regime: $\Delta>10^{-2}$ is represented by light pink color, where $\mathcal{F}_{\mu\mu}\sim L^{2}$. The parameter $\mu=2$ is fixed throughout.
II Model
Model.– We consider the one-dimensional Kitaev lattice model [47, 48], $\hat{H}=\hat{H}_{1}+\hat{H}_{\Delta}$, where
@@ II Model
Here $\hat{c}^{\dagger}_{j}$ ($\hat{c}_{j}$) is the creation (annihilation) operator at site $i$ of a spinless fermion, $L$ is the number of lattice sites, $\mu$ is an on-site potential and $\Delta$ is the superconducting $p$-wave pairing amplitude. This model has a rich topological phase diagram [49, 50, 51, 52], which can be characterized via a topological invariant, the winding number, $w$. Two critical lines $\mu=\pm 2$ and $\Delta=0$ set apart three topologically distinct phases: a trivial phase for $|\mu|>2$ with $w=0$, one topologically non-trivial phase for $|\mu|<2$ with $w=1$ for $\Delta>0$, and another topologically non-trivial phase for $|\mu|<2$ with $w=-1$ for $\Delta<0$. The one-dimensional Kitaev model hosts Majorana Zero Modes [53, 54] in the topological phases at both of its ends. The intersections of two critical lines marks two multicritical points in the system. The multicritical point connects two topologically non-trivial phases with a topologically trivial phase. Topological properties of a $p$-wave spin-less superconductor have been studied from different points of views in recent years [55, 56, 57, 58, 59, 60, 51, 61]. There has been several proposals and experimental efforts for physical realization of the Kitaev $p$-wave model [62, 56, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72].
@@
III Multi-Parameter sensing
Essential theory of single- and multi-parameter sensing.– Recently, sensing for more than one parameter or multi-parameter sensing has gained interest [73, 74, 75]. The parameters $\mu$, $\Delta$ of the Hamiltonian, $\hat{H}$, are encoded in the ground state. These parameters are changed, which gets reflected in the ground state. By employing the methods of the quantum Fisher information matrix (QFIM), we get a $2\times 2$ matrix $\mathcal{F}$ given by $\mathcal{F}_{ab}=4\text{Re}(\langle\partial_{a}\Psi|\partial_{b}\Psi\rangle-\langle\partial_{a}\Psi|\Psi\rangle\langle\Psi|\partial_{b}\Psi\rangle)$, where $a,b=\mu,\Delta$ and $|\Psi\rangle$ is the probe state. While considering multi-parameter sensing, the equally weighted uncertainty is given by,
$\delta\mu^{2}+\delta\Delta^{2}\geq\frac{1}{m}\mathcal{G}^{-1},$ (3)
where $\mathcal{G}:=(\text{Tr}[\mathcal{F}^{-1}])^{-1}=\frac{\mathcal{F}_{\mu\mu}\mathcal{F}_{\Delta\Delta}-\mathcal{F}_{\Delta\mu}\mathcal{F}_{\mu\Delta}}{\mathcal{F}_{\mu\mu}+\mathcal{F}_{\Delta\Delta}}$ and $m$ is the number of repetition of the sensing protocol. The quantity $\mathcal{G}^{-1}$ is the total uncertainty for the equally weighted multi-parameter estimation. For the single-parameter estimation case, i.e., when one of the two parameters is precisely known, it is sufficient to compute the relevant diagonal element of the QFIM. Further discussion on this is presented in Appendix A.
Symmetry-breaking assisted parameter estimation at multicriticality.– In absence of the pairing term, i.e., for $\Delta=0$, the system has the global $U(1)$ symmetry as the Hamiltonian is invariant under the transformation $\hat{c}_{i}\to e^{i\phi}\hat{c}_{i}$, where $\phi$ is a site-independent phase. This global symmetry implies the conservation of the total fermion number. The corresponding system is gapless for parameter regime $|\mu|\leq 2$. Introduction of $\hat{H}_{\Delta}$ breaks the $U(1)$ symmetry, and a gap opening is occurred. The phase diagram is gaped everywhere in the $\mu-\Delta$ plane except at the critical lines at $|\mu|=2$. The basic idea of the symmetry-breaking quantum enhancement of precision is the following: For parameter estimation of $\Delta$, assuming $\mu$ as a known parameter that can be precisely tuned, the many-body ground state of the system is prepared near the multi-critical point ($\mu=2$ and $\Delta\to 0$). If the system is subjected to a small variation in $\Delta$, it is bound to happen along the gapless critical line at $\mu=2$. As we show below, this gives rise to Heisenberg scaling in $\mathcal{F}_{\Delta\Delta}$ – thanks to the $U(1)$ symmetry-breaking. This is not the case with the parameter estimation of $\mathcal{F}_{\mu\mu}$. It vanishes on the gapless critical line, and must involve a gap-to-gapless transition by setting $\Delta$ non-zero. Details related to this will be furnished later. The idea, however, can be generalized for simultaneous estimation of two-parameters by combing the effects of, both, symmetry-breaking and a gap-to-gapless transition. A super-Heisenberg scaling is obtained under such scenario in $\mathcal{G}$. The is the main idea that we pitch in this work and is schematically presented in Fig. 1. In the following we present the results.
QFIM near criticality: Leading-order scaling.– We begin with summarizing the key results before we proceed to provide further details. We use fidelity based definition of $\mathcal{F}_{ab}$ for deriving the elements of QFIM for the ground state $|\Phi_{G}\rangle$ (Eq. (2)). Following expressions are obtained: $\mathcal{F}_{\mu\mu}=\sum_{k}\frac{\Delta^{2}\sin^{2}k}{\mathcal{E}_{k}^{4}}$, $\mathcal{F}_{\mu\Delta}=\sum_{k}\frac{\Delta z_{k}\sin^{2}k}{\mathcal{E}_{k}^{4}}$ and $\mathcal{F}_{\Delta\Delta}=\sum_{k}\frac{z_{k}^{2}\sin^{2}k}{\mathcal{E}_{k}^{4}}$. Unlike classical Fisher Information, which is additive, Quantum Fisher Information is super-additive. Hence, what are of prime interests are the system-size scalings of the elements of QFIM and the ultimate precision estimator $\mathcal{G}$. By calculating the limits of $\mathcal{F}_{\mu\mu}$ as $\Delta\to 0$ and $\mu\to 2$, we obtain closed form expressions of the leading order scalings as, $\mathcal{F}_{\mu\mu}\sim{\Delta^{2}L^{6}}/\pi^{6}$, $\mathcal{F}_{\Delta\Delta}\sim L^{2}/8$, $\mathcal{F}_{\mu\Delta}\sim{\Delta L^{4}}/{\pi^{4}}$ and $\mathcal{G}\sim{(\pi^{2}-8)\Delta^{2}L^{6}}/{\pi^{8}}$. Thus, we observe super-Heisenberg scaling in all quantities of our interest except $\mathcal{F}_{\Delta\Delta}$ for which it saturates the Heisenberg limit. These scalings of the QFIM are remarkable results establishing the importance of the multicritical point and symmetry-breaking for precision measurement of multiple parameters. This constitutes the main result of our work. The details of the derivations for obtaining the analytical expressions are presented in Appendix B.
Figure 3: Variation of $\mathcal{F}_{\Delta\Delta}$ and $\mathcal{G}$ near the multicritical point: (a) $\mathcal{F}_{\Delta\Delta}$ for system sizes $L$ = 400 (solid blue line), 600 (dot-dashed orange line), 800 (dotted green line), 1000 (dashed red line) are presented with respect to $\Delta$ keeping a fixed $\mu=2.0$. This is a representative case, we observe similar nature for various other system-sizes as well. The finite-size scaling of $\mathcal{F}_{\Delta\Delta}$ and $\mathcal{G}$ are presented in (b) and (c), respectively. Setting $\mu=2$, $\mathcal{F}_{\Delta\Delta}$ and $\mathcal{G}$ at $\Delta=10^{-7}$ (circular markers) scale as $\mathcal{F}_{\Delta\Delta}^{a}\sim L^{2}$ and $\mathcal{G}^{a}\sim L^{6}$, while at $\Delta=0.7$ (square markers) they scale as $\mathcal{F}_{\Delta\Delta}^{b}\sim L$ and $\mathcal{G}^{b}\sim L$. The straight lines are the best fit. (d) We present the scaling exponents $\beta$ of $\mathcal{F}_{\Delta\Delta}$ and $\mathcal{G}$ against the parameter $\Delta$. We identify three regions, (i) low $\Delta$ regime: $\Delta<10^{-5}$ is represented by light blue color where $\mathcal{F}_{\Delta\Delta}\sim L^{2}$ and $\mathcal{G}\sim L^{6}$; (ii) intermediate $\Delta$ regime : $10^{-5}\leq\Delta\leq 0.03$ represented by light green color where $\beta$ transitions for both $\mathcal{F}_{\Delta\Delta}$ and $\mathcal{G}$; (iii) large $\Delta$ regime: $\Delta>0.03$ is represented by light pink color where $\mathcal{F}_{\Delta\Delta},\mathcal{G}\sim L$ The parameter $\mu=2$ is fixed throughout.
Figure 4: Variation of $\mathcal{G}$ near multicriticality : The quantity $\mathcal{G}$ is presented with respect to the tuning parameters $\mu$ and $\Delta$ in the vicinity of the multicriticality. The probe considered is the ground state of the Kitaev model on a one-dimensional lattice with system size of $L=1000$.
$\mathcal{F}_{\mu\mu}$ at multicriticality and beyond.– As mentioned before, $\mathcal{F}_{\mu\mu}=\sum_{k}\frac{\Delta^{2}\sin^{2}k}{\mathcal{E}_{k}^{4}}$. Evidently, $\mathcal{F}_{\mu\mu}$ vanishes on the critical line, $\Delta=0$. A non-vanishing value $\mathcal{F}_{\mu\mu}$ requires to prepare the system at a finite $\Delta$. In Fig. 2(a), we show the nature of $\mathcal{F}_{\mu\mu}$ with respect to $\mu$, for a representative case, where $\Delta$ is set at 0.001. $\mathcal{F}_{\mu\mu}$ develops a peak at $\mu\sim 2.0$. The peak corresponds to a gap-to-gapless transition. In Fig. 2(b), we present the finite-size scaling of $\mathcal{F}_{\mu\mu}$ at $\mu=2$ and for two representative cases of $\Delta=10^{-7}$ with scaling exponent $\beta=6$ and $\Delta=0.7$ with $\beta=1$. This change in scaling exponent is gradual with change in the value of $\Delta$ and is presented in Fig. 2(c). We observe that based on the value of $\beta$, the parameter range of $\Delta$ can be divided into three parts: small $\Delta$-regime with super-Heisenberg $\beta=6$, intermediate $\Delta$-regime where the $\beta$ transitions from super-HL to SQL and large $\Delta$-regime where we observe SQL scaling, i.e., $\beta=1$.
$\mathcal{F}_{\Delta\Delta}$ and $\mathcal{G}$ at multicriticality and beyond.– Following expressions are obtained: $\mathcal{F}_{\Delta\Delta}=\sum_{k}\frac{z_{k}^{2}\sin^{2}k}{\mathcal{E}_{k}^{4}}$ and $\mathcal{F}_{\mu\Delta}=\sum_{k}\frac{\Delta z_{k}\sin^{2}k}{\mathcal{E}_{k}^{4}}$. $\mathcal{G}$ can be computed from there-off as it is a function of $\mathcal{F}_{ab}$, where $a,b=\mu,\Delta$. Validity of these expressions are further confirmed numerically. It is, however, imperative that while changing both the parameters simultaneously in an adiabatic manner, it is done in such a fashion that the parameter $\mu$ does not follow the critical $\Delta=0$, along which $\mathcal{F}_{\mu\Delta}$ vanishes. Hence, the variation in $\mu$ must be guided entirely within the symmetry-broken phase with non-vanishing $\Delta$, and such that it encounter a gap-to-gapless transition. The parameter $\Delta$, however, can be navigated along the critical line, $\mu=2$, such that it involves the symmetry-breaking transition. The combined effects of the gap-to-gapless transition and the symmetry-breaking gives rise to a super-Heisenberg scaling, as we exemplify below. It is by now obvious that the case of single parameter estimation $\mathcal{F}_{\Delta\Delta}$ can be done solely via symmetry-breaking by tuning $\mu$ precisely at $\mu=2$ and by preparing the system at or near the multicritical point.
We present a set of results displaying the behavior and scaling of $\mathcal{F}_{\Delta\Delta}$ and $\mathcal{G}$ with respect to the $\Delta$ in Fig. 3(a-d). In Fig. 3(a) we present how $\mathcal{F}_{\Delta\Delta}$ varies with the parameter $\Delta$ when $\mu=2$ for various system-sizes. We observe that $\mathcal{F}_{\Delta\Delta}$ saturates to a particular maximum value as $\Delta\rightarrow 0$. The quantity $\mathcal{F}_{\Delta\Delta}$ at $\mu=2$ becomes independent of $\Delta$ ($\mathcal{F}_{\Delta\Delta}\sim L^{2}/8$) for very small values of $\Delta$, say up-to $\Delta^{*}$. As the value of $\Delta$ increases, $\mathcal{F}_{\Delta\Delta}$ decreases gradually with increase in $\Delta$. It can be expected that $\Delta^{*}\to 0$ in the thermodynamic limit. In Fig. 3(b) the finite-size scaling of $\mathcal{F}_{\Delta\Delta}$ is presented at $\mu=2$ for two representative cases, $\Delta=10^{-7}$ with super-HL scaling $\beta=2$ and $\Delta=0.7$ with SQL scaling $\beta=1$. We provide a similar analysis for $\mathcal{G}$ in Fig. 3(c), where we show that $\mathcal{G}$ has super-HL of $\beta=6$ at $\Delta=10^{-7}$ but a SQL scaling at a larger $\Delta=0.7$. In Fig. 3(d) we provide the change in the scaling exponent of $\mathcal{F}_{\Delta\Delta}$ and $\mathcal{G}$ with change in $\Delta$. The range of values of $\Delta$ can be divided into three subparts: small $\Delta$ region where the scaling of $\mathcal{F}_{\Delta\Delta}$ saturates to HL $\beta=2$ as $\Delta$ decreases and $\mathcal{G}$ has a super-HL scaling of 6. In the intermediate $\Delta$ range the scaling exponents of both of the quantities transition to SQL vale of $\beta=1$. In the large $\Delta$ region the scaling exponents of both the quantities saturate to 1.
In Fig. 4, we present the variation of $\mathcal{G}$ with respect to the parameters $\mu$ and $\Delta$ for system-size $L=1000$. This is a representative case, and we observe similar behavior for other system-sizes as well. We observe that the value of $\mathcal{G}$ is maximum around the critical point of $\mu=2.0$. The nature of $\mathcal{G}$ is similar to that of $\mathcal{F}_{\Delta\Delta}$ in the topological regime ($\mu<2.0$) and is similar to $\mathcal{F}_{\mu\mu}$ in the trivial phase ($\mu>2.0$). The scaling exponent $\beta$ of $\mathcal{G}$ saturates to 6 for smaller values of $\Delta$ and saturates to 1 for larger values. The precision limit is inversely proportional to the quantity $\mathcal{G}$. Hence, total uncertainty for equally weighted multi-parameter estimation is $\delta\mu^{2}+\delta\Delta^{2}\propto\mathcal{G}^{-1}\sim L^{-\beta}$, where $\beta$ is the scaling exponent of $\mathcal{G}$, viz. $\mathcal{G}\sim L^{\beta}$. Thus, larger the scaling of $\mathcal{G}$, the more precisely can one simultaneously sense $\mu$ and $\Delta$ for larger system-sizes. Hence, the ground state of the Kitaev model on a one-dimensional lattice, when used as a probe, is a resource for obtaining quantum-enhanced multi-parameter sensitivity.
IV Discussions
There has been several proposals for engineering many-body quantum sensors by exploiting criticality. In all these works reported so far, gap-to-gapless transition has been identified as the key resource for achieving the quantum enhanced sensing. Other important concept associated with quantum criticality, such as symmetry-breaking, has not been shown so far to play any crucial role in parameter estimation. In this work we demonstrate for the first time that symmetry-breaking can also be a useful resource in parameter estimation. We show this by exploiting the the ground-state of the Kitaev model on a one-dimensional lattice as a candidate for a multi-parameter sensor.
The probe acting as an adiabatic sensor provides quantum-enhanced sensitivity near the multicritical point. The system admits a topological phase for $|\mu|\leq 2$ and a trivial phase beyond that. The topological phase is characterized by a non-zero winding number. In addition, the $\Delta=0$ point brings about a flip in the sign of the winding number when the system is in the topological phase. At the critical point of $\mu=2.0$ and $\Delta=10^{-7}$, we obtain single-parameter scaling of the QFI-s $\mathcal{F}_{\mu\mu}\sim L^{6}$ and $\mathcal{F}_{\Delta\Delta}\sim L^{2}$, whereas the multi-parameter sensitivity at this point also touches $\mathcal{G}\sim L^{6}$ mark. Beyond this point, we do get quantum advantage in the form of super-SQL scaling for an extended region of the parameter-space. In our work, we present a phase diagram of the system and discuss how the QFIM and total uncertainty limit behave in the parameter-space of either side of the critical line.
In summary, thus, it can be said that the symmetry-breaking transition can be used as a resource, similar to gap-gapless transition, for engineering a new class of quantum sensors, as demonstrated by considering the topological Kitaev chain. It will be interesting to pursue further investigations in future, such as understanding the effects of long-range tunneling or long-range pairing on the sensing protocols proposed in this work.
Appendix A Theory of multi-parameter Sensing
[…]

### v2

@@ Abstract
Quantum criticality has been demonstrated as a useful quantum resource for parameter estimation. This includes second-order, topological and localization transitions. In all these works reported so far, gap-to-gapless transition at criticality has been identified as a crucial resource for achieving the quantum-enhanced sensing, although there are several important concepts associated with criticality, such as long-range correlation, symmetry breaking. In this work, we show that symmetry-breaking alone can drive a quantum-enhanced sensing, even without any gap-to-gapless transition. We analytically demonstrate that the estimation of the superconducting pairing amplitude in the one-dimensional Kitaev model achieves Heisenberg scaling when the system is prepared near a multicritical point and is varied along a gapless critical line, implying symmetry breaking as a standalone metrological resource. Extending our analysis in the realm of simultaneous multiparameter estimation of both the pairing term and the chemical potential, we show that it is possible to obtain $L^{6}$ scaling in a narrow parameter range, but with definite observable consequence, where the quantum advantage is assisted by gap-to-gapless transition as well. Our work thus identifies a new resource for criticality-enhanced quantum sensing, and also suggests multicritical systems as useful platform for multiparameter sensing.
@@ I Introduction
The quantum Cramér-Rao bound provides the ultimate limit of precision within which an unknown parameter can be estimated in the quantum systems [1, 2, 3, 4]. The uncertainty $\epsilon_{q}$ in parameter estimation is related to a very special quantity, quantum Fisher information (QFI). The bound suggests that $\epsilon_{q}\geq(MF_{Q})^{-1}$ for the quantum systems, where $F_{Q}$ is the QFI and $M$ is the number of repetition of the sensing protocol. Quantum systems might provide a true quantum edge over classical systems when it comes to accurately measuring an unknown variable. This is demonstrated by how the uncertainty scales with the size of the probe. Given $\epsilon_{q}\propto L^{-\nu}$, where $L$ is the probe size and $\nu$ is the associate scaling exponent, the classical probes can at-best scale linearly with system size, i.e., $\nu=1$, which corresponds to the standard quantum limit (SQL). However, it is possible to exploit quantum phenomena, such as quantum correlation or cooperative phenomena, for achieving quantum-enhanced sensitivity, where $\nu>1$. The so-called Heisenberg limit (HL) corresponds to $\nu=2$ [5]. Subsequently, quantum sensing [6, 2, 7] has taken the centre stage of quantum technology with several applications, and has been investigated from different points of views [5, 8, 9, 10, 11, 12, 13].
The cooperative quantum phenomena, such as quantum phase transitions in quantum many-body (QMB) systems, has turned out to be an extremely useful quantum resource for engineering new types of quantum sensors. In literature, quantum criticality has been used to undertake various metrological tasks [14, 15, 16, 17]. Quantum criticality, in its various forms, have been utilized to attain HL limit, or for even realizing beyond HL limits. This includes symmetry-breaking second-order quantum phase transitions [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 22, 39, 40, 41], as well as symmetry-protected quantum phase transitions, such as localization transition [42, 43, 44, 45] and topological phase transition (TPT) [46, 46, 47, 48, 49, 50, 51], and non-Hermitian topological systems [52, 53, 54, 55]. It has been demonstrated in literature that when considering other resources like preparation time, the apparent super-HL scaling can diminish to HL [25, 7]. However, recently it has been shown that even after considering preparation time, super-HL scaling can be preserved in certain systems [42]. The effectiveness of super-HL scaling quantum sensing based on QFI and quantum Cramér-Rao bound has been studied extensively in the literature and it has been found that sub-HL strategies are ineffective [56, 57].
@@
There are two different key motivations for probing different kinds of quantum criticality - first, it is important to understand the basic ingredients leading to quantum-enhanced sensitivity, e.g. the role of gap closing or symmetry-breaking; secondly, there is a constant quest for identifying experimentally realizable QMB systems with super-Heisenberg scaling property characterized by large scaling exponent. It has been shown that the gap closing at criticality leads to a scaling of $F_{Q}\sim L^{(2/\nu)}$, where $L$ is the system size, and the critical exponent $\nu$ characterizing the divergence of the correlation length (localization length) near the criticality in case of the second-order quantum phase transition (localization transition). Although there are many key aspects associated with quantum criticality, such as symmetry breaking, gap closing, long-range correlation, only the gap closing has been identified as the crucial ingredient for obtaining quantum-enhanced sensitivity, implying $\nu<2$. It remains unclear whether other critical properties, such as symmetry breaking, can independently provide quantum advantage.
In this work we show that gap-to-gapless transition is not always required for quantum-enhanced sensing. We show that symmetry breaking alone, when the system is prepared near a multicritical point, can be sufficient to obtain Heisenberg-limited scaling in single-parameter estimation. To the best of our knowledge, such symmetry-breaking-assisted quantum advantage for quantum sensing has not been reported in literature previously. We establish this result while performing single-parameter estimation of the superconducting pairing gap in the one-dimensional Kitaev model. We then extend the analysis to simultaneous multiparameter estimation of the pairing term and the onsite potential, where multicriticality allows super-Heisenberg scaling. The individual contributions of symmetry breaking and gap closing, however, can no longer be separated, as both effects coexist.
The rest of the paper is organized as follows. In Sec. II, we provide a brief overview of the Kitaev model on one-dimensional lattice and briefly discuss the theory of multi-parameter sensing. In Sec. III and IV, we consider single-parameter sensing. There in Sec. III we demonstrate how symmetry-breaking can be exploited as a resource for quantum sensing of the pairing amplitude and establish one of the central results of this work. In Sec. V, the QFI matrix for multiparameter sensing is presented. We demonstrate that multicritical point can be exploited for quantum enhanced sensing of multiple parameters, representing another main result of this work. In Sec. VI, we study the parameter regime in which the proposed sensor provides quantum advantage. We also examine how the sensor works, when state preparation time is taken into consideration. Finally, we present our concluding remarks and discussions in Sec. VII.
II Preliminaries
II.1 Model
We consider the one-dimensional Kitaev lattice model [58, 59], $\hat{H}=\hat{H}_{1}+\hat{H}_{\Delta}$, where
@@ II.1 Model
Here $\hat{c}^{\dagger}_{j}$ ($\hat{c}_{j}$) is the creation (annihilation) operator at site $i$ of a spinless fermion, $L$ is the number of lattice sites, $\mu$ is an on-site potential and $\Delta$ is the superconducting $p$-wave pairing amplitude. This model has a rich topological phase diagram [60, 61, 62, 63], which can be characterized via a topological invariant, the winding number, $w$. Two critical lines $\mu=\pm 2$ and $\Delta=0$ set apart three topologically distinct phases: a trivial phase for $|\mu|>2$ with $w=0$, one topologically non-trivial phase for $|\mu|<2$ with $w=1$ for $\Delta>0$, and another topologically non-trivial phase for $|\mu|<2$ with w=1w=-1 for $\Delta<0$. A schematic representation of these phases and the associated critical lines is shown in Fig. 1. The one-dimensional Kitaev model hosts Majorana Zero Modes [64, 65] in the topological phases at both of its ends. The intersections of two critical lines mark two multicritical points in the system. The multicritical point connects two topologically non-trivial phases with a topologically trivial phase. Topological properties of a $p$-wave spin-less superconductor have been studied from different points of views in recent years [66, 67, 68, 69, 70, 71, 62, 72]. There has been several proposals and experimental efforts for physical realization of the Kitaev $p$-wave model [73, 67, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83].
@@
Figure 1: Schematic for quantum-enhanced sensing assisted by symmetry-breaking: The phase diagram shows different topological phases in the $\mu$-$\Delta$ plane. The red and the green lines are gapless. In the left panel, the protocol of sensing $\Delta$ is presented. $\mathcal{F}_{\Delta\Delta}$ has an quantum-enhanced scaling when the system is prepared at the critical line of $\mu=\mu_{c}$ (near the multicritical point, at A), and driven to, say B. We note that this is enhancement of scaling is not because of gap-to-gapless transition, but rather due to breaking of $U(1)$ symmetry. In the multi-parameter regime, one can envision a case where system is prepared at C, and both the parameters are maneuvered along the directed arrows. Since, we are in a two-dimensional plane, the point C can be approached from multiple directions. This leads to a quantum-enhanced scaling for the multi-parameter precision estimator.This involves, both, a gap-to-gapless transition and the symmetry-breaking.
II.2 Theory of single- and multi-parameter sensing
Let us consider a set of parameters, $\mathbf{x}=\{x_{a}\}_{a=1}^{n}$, encoded in some density matrix $\rho(\mathbf{x})$. The precision limits of any parameter that is being sensed is given by a $n\times n$ matrix called the quantum Fisher Information matrix (QFIM). The elements of the QFIM represented by $\mathcal{F}$ is given by, $\mathcal{F}_{ab}:=\frac{1}{2}\text{Tr}\left(\rho\{L_{a},L_{b}\}\right)$, where $\{\cdot,\cdot\}$ represents the anti-commutation and the operators $L_{a}$ and $L_{b}$ are the symmetric logarithmic derivatives (SLD) with respect to $x_{a}$ and $x_{b}$, respectively. This is given by the equation, $\partial_{i}\rho=\frac{1}{2}\left(\rho L_{i}+L_{i}\rho\right)$, where, $i=a,b$ and $\partial_{i}$ is the partial derivative with respect to $x_{i}$ viz. $\partial_{i}=\partial/\partial x_{i}$. In this work, we specifically work with the many-body ground state of the Kitaev model. Hence our state of interest is a pure state. For pure states, the elements of QFIM reduces to $\mathcal{F}_{ab}=4\text{Re}\left(\langle\partial_{a}\psi|\partial_{b}\psi\rangle-\langle\partial_{a}\psi|\psi\rangle\langle\psi|\partial_{b}\psi\rangle\right)$. Here the state $|\psi\rangle=|\psi(\mathbf{x})\rangle$ encodes the parameters $\mathbf{x}$. It can be seen that the diagonal elements of QFIM - $\mathcal{F}_{aa}$ reduces to the QFI for parameter $x_{a}$.